d) The table below represents choices and rewards available to a student attending BECO 101 for her use of 1600hrs to 1800hrs Mondays’ class time. S/No Choice Reward 1. Attend class 18,210 2. Sleep 19,480 3. Watch a movie 21,494 4. Chat with friends 19,901 5. Cook Dinner 27,483 Required: i. Generate a column of the opportunity costs associated with each choice. (5 marks) ii. Mindful of the Rationality assumption in economics, what choice would you advise the student to take? Give reasons. (5 marks)
1
Expert's answer
2021-11-01T12:48:16-0400

a.Opportunity cost=foregone alternative"\div" selected alternative

"\frac{210}{18}=15.56\\
\frac{480}{19}=25.26\\
\frac{494}{21}=23.52\\
\frac{901}{19}=47.42\\
\frac{483}{17}=28.41"


b.choice 5 because it has high opportunty cost